# Anomalous proximity effect under Andreev and Majorana bound states

### v2

@@ 3.2 Dependence on S Region Length
Figure 7: LDOS for long S region ($L_{\rm S}=100a$) junctions evaluated at the leftmost site of the N region ($i=1$) as a function of energy $E$ and scalar disorder strength $W$ with $L_{\rm DN}=2a$ (a,f), $L_{\rm DN}=20a$ (b,g), and $L_{\rm DN}=40a$ (c,h). Panels (d,i) show line cuts of the LDOS in (a,f,b,g) for $W/\mu_{S}=0,1,2$. Panels (e,j) show the spatial profile of the LDOS at $E\approx 0$ for $W/\mu_{S}=0,1,2$ with $L_{\rm DN}=20a$. Top row (a,b,c,d,e): trivial phase ($B=0.825B_{c}$). Bottom row (f,g,h,i,j): topological phase ($B=1.2B_{c}$). Other parameters used are: $L_{\rm S}=100a$, $L_{\rm N}=40a$, $a=50$ nm, $\mu_{S}=0.5$ meV, $\mu_{N}=0.1$ meV, $\Delta=0.25$ meV, $\alpha=20$ meV nm, $t=1$ meV, and $\eta=0.001$ meV.
Figure 8: Absolute value of the spin-triplet pair correlations $|\vec{d}|$ for long S region ($L_{\rm S}=100a$) junctions evaluated at the leftmost site of the N region ($i=1$) as a function of energy $E$ and scalar disorder strength $W$ with $L_{\rm DN}=2a$ (a,f), $L_{\rm DN}=20a$ (b,g), and $L_{\rm DN}=40a$ (c,h). Panels (d,i) show line cuts of the spin-triplet pair correlations in (a,f,b,g) for $W/\mu_{S}=0,1,2$. Panels (e,j) show the spatial profile of the spin-triplet pair correlations at $E\approx 0$ for $W/\mu_{S}=0,1,2$ with $L_{\rm DN}=20a$. Top row (a,b,c,d,e): trivial phase ($B=0.825B_{c}$). Bottom row (f,g,h,i,j): topological phase ($B=1.2B_{c}$). Other parameters used are: $L_{\rm S}=100a$, $L_{\rm N}=40a$, $a=50$ nm, $\mu_{S}=0.5$ meV, $\mu_{N}=0.1$ meV, $\Delta=0.25$ meV, $\alpha=20$ meV nm, $t=1$ meV, and $\eta=0.001$ meV.
Figure 9: Absolute value of the spin-singlet pair correlation for long S region ($L_{\rm S}=100a$) junctions evaluated at the leftmost site of the N region ($i=1$) as a function of energy $E$ and scalar disorder strength $W$ with $L_{\rm DN}=2a$ (a,f), $L_{\rm DN}=20a$ (b,g), and $L_{\rm DN}=40a$ (c,h). Panels (d,i) show line cuts of the spin-singlet pair correlations in (a,f,b,g) for $W/\mu_{S}=0,1,2$. Panels (e,j) show the spatial profile of the spin-singlet pair correlations at $E\approx 0$ for $W/\mu_{S}=0,1,2$ with $L_{\rm DN}=20a$. Top row (a,b,c,d,e): trivial phase ($B=0.825B_{c}$). Bottom row (f,g,h,i,j): topological phase ($B=1.2B_{c}$). Other parameters used are: $L_{\rm S}=100a$, $L_{\rm N}=40a$, $a=50$ nm, $\mu_{S}=0.5$ meV, $\mu_{N}=0.1$ meV, $\Delta=0.25$ meV, $\alpha=20$ meV nm, $t=1$ meV, and $\eta=0.001$ meV.
@@ 4.1 Anomalous Proximity Effect for Short S Regions
To further understand the nature of the ZEP and the features observed in the LDOS for the short superconducting segment, we now turn our attention to the the proximity-induced spin-triplet and spin-singlet pair correlations. In Fig. 5, we show the absolute value of the spin-triplet pair correlations $|\vec{d}|$ at the leftmost site of the N region ($i=1$) as a function of energy $E$ and disorder strength $W$ for different lengths of the DN region ($L_{\rm DN}$). In the trivial phase (top row of Fig. 5), we observe a ZEP in $|\vec{d}|$ in the clean limit. As disorder is introduced, the peak splits rapidly, evolving into a fan-shaped structure similar to that observed in the LDOS. Taking line cuts at different values of $W$, we clearly see the flattening of the ZEP into almost a plateau as a function of energy. The spatial profile of the spin-triplet pairing at $E\approx 0$ (Fig. 5(e) shows that in the clean limit (grey curve), triplet correlations exhibit a large value at the DN/S interface where the ABS is localized and it penetrates significantly into the N region. However, with increasing disorder (red and cyan curves), the magnitude of these zero-energy triplet correlations is substantially reduced throughout the entire N region, and their characteristic oscillatory pattern is dampened, especially within the DN segment.
@@ Acknowledgements.
We thank S. Ikegaya for fruitful discussions. E. A. acknowledges financial support from Nagoya University and Mitsubishi Foundation. Y. T. acknowledges support from JSPS with Grants-in-Aid for Scientific Research (KAKENHI Grant Nos. 23K17668,24K00583,24K00578,24K00556,25H00609, and 25H00613). J. C. acknowledges financial support from the Carl Trygger’s Foundation (Grant No. 22: 2093), the Sweden-Japan Foundation (Grant No. BA24-0003), and the Swedish Research Council (Vetenskapsrådet Grant No. 2021-04121).

### v3

@@ 3.2 Dependence on S Region Length
Figure 7: Ensemble-averaged LDOS for long S region ($L_{\rm S}=100a$) junctions evaluated at the leftmost site of the N region ($i=1$) as a function of energy $E$ and scalar disorder strength $W$ with $L_{\rm DN}=2a$ (a,f), $L_{\rm DN}=20a$ (b,g), and $L_{\rm DN}=40a$ (c,h). Panels (d,i) show line cuts of the LDOS in (a,f,b,g) for $W/\mu_{S}=0,1,2$. Panels (e,j) show the spatial profile of the LDOS at $E\approx 0$ for $W/\mu_{S}=0,1,2$ with $L_{\rm DN}=20a$. Top row (a,b,c,d,e): trivial phase ($B=0.825B_{c}$). Bottom row (f,g,h,i,j): topological phase ($B=1.2B_{c}$). Other parameters used are: $L_{\rm S}=100a$, $L_{\rm N}=40a$, $a=50$ nm, $\mu_{S}=0.5$ meV, $\mu_{N}=0.1$ meV, $\Delta=0.25$ meV, $\alpha=20$ meV nm, $t=1$ meV, and $\eta=0.001$ meV.
Figure 8: Ensemble-averaged Absolute value of the spin-triplet pair correlations $|\vec{d}|$ for long S region ($L_{\rm S}=100a$) junctions evaluated at the leftmost site of the N region ($i=1$) as a function of energy $E$ and scalar disorder strength $W$ with $L_{\rm DN}=2a$ (a,f), $L_{\rm DN}=20a$ (b,g), and $L_{\rm DN}=40a$ (c,h). Panels (d,i) show line cuts of the spin-triplet pair correlations in (a,f,b,g) for $W/\mu_{S}=0,1,2$. Panels (e,j) show the spatial profile of the spin-triplet pair correlations at $E\approx 0$ for $W/\mu_{S}=0,1,2$ with $L_{\rm DN}=20a$. Top row (a,b,c,d,e): trivial phase ($B=0.825B_{c}$). Bottom row (f,g,h,i,j): topological phase ($B=1.2B_{c}$). Other parameters used are: $L_{\rm S}=100a$, $L_{\rm N}=40a$, $a=50$ nm, $\mu_{S}=0.5$ meV, $\mu_{N}=0.1$ meV, $\Delta=0.25$ meV, $\alpha=20$ meV nm, $t=1$ meV, and $\eta=0.001$ meV.
Figure 9: Ensemble-averaged Absolute value of the spin-singlet pair correlation for long S region ($L_{\rm S}=100a$) junctions evaluated at the leftmost site of the N region ($i=1$) as a function of energy $E$ and scalar disorder strength $W$ with $L_{\rm DN}=2a$ (a,f), $L_{\rm DN}=20a$ (b,g), and $L_{\rm DN}=40a$ (c,h). Panels (d,i) show line cuts of the spin-singlet pair correlations in (a,f,b,g) for $W/\mu_{S}=0,1,2$. Panels (e,j) show the spatial profile of the spin-singlet pair correlations at $E\approx 0$ for $W/\mu_{S}=0,1,2$ with $L_{\rm DN}=20a$. Top row (a,b,c,d,e): trivial phase ($B=0.825B_{c}$). Bottom row (f,g,h,i,j): topological phase ($B=1.2B_{c}$). Other parameters used are: $L_{\rm S}=100a$, $L_{\rm N}=40a$, $a=50$ nm, $\mu_{S}=0.5$ meV, $\mu_{N}=0.1$ meV, $\Delta=0.25$ meV, $\alpha=20$ meV nm, $t=1$ meV, and $\eta=0.001$ meV.
@@ 4 LDOS and proximity-induced pair correlations
4.1 Physical picture: Disorder and the role of chiral symmetry
Before we present our numerical results, it is useful to establish the physical mechanism governing the stability of the ZEP against disorder. The key to this protection is the chiral symmetry of the BdG Hamiltonian PhysRevB.94.054512 (66, 72, 73). Generally speaking, for any chiral symmetric systems, any zero-energy state (ZES) is necessarily an eigenstate of the chiral operator $\Gamma$, and has a definite chirality ($\gamma=\pm 1$). Meanwhile, any nonzero energy state must be an equal superposition of states with opposite chiralities. This is a direct consequence of the the chiral symmetry condition ($\Gamma H\Gamma=-H$). If chiral symmetry preserving disorder (e.g. scalar disorder) is introduced into the system, disorder can only move a ZES to a finite energy if it can couple it with another ZES of the opposite chirality. The number of robust ZES is therefore given by the index $N_{ZES}=|N_{+}-N_{-}|$, which represents the imbalance between the number of states with positive ($N_{+}$) and negative ($N_{-}$) chirality PhysRevB.94.054512 (66).
This framework provides a clear, microscopic picture for our results that hinges on the spatial properties of the chiral partner states. For the topological MBSs, the two states with opposite chirality are the two MBSs themselves, localized at opposite ends of the superconducting region. In a long S-region, these two states are spatially well-separated. A local perturbation like scalar disorder in the DN region cannot simultaneously couple these distant states. Therefore, the MBS at the DN/S interface remains robustly pinned at zero energy. The MBSs in a short S-region are fragile for the same reason. Mainly, their wavefunctions overlap significantly, allowing local disorder to couple them and left them away from zero energy.
For the trivial ABS, the situation is drastically different. The ABS itself is a superposition of two spatially overlapping eigenstates of the chiral operator with opposite chirality. One eigenstate is localized precisely at the DN/S interface and is responsible for the localized behavior observed in the top panels of Fig. 2(c,f,i), while the other is a more delocalized, particle-in-a-box-like state within the normal region and is responsible for the oscillatory behavior of the wavefunction in the top panels of Fig. 2(c,f,i). Because these two opposite-chirality components coexist in the same location, local disorder can easily couple them, lifting the degeneracy and pushing the ABS to a finite energy. This explains its fragility.
Crucially, the ZEP observed in the LDOS probed at the leftmost site of the junction should reflect the physics discussed above. In particular, we should expect fragile ZEPs in the trivial phase and in the topological phase of junctions with short S region. In contrast, the ZEP should be robust in the topological phase when the MBSs are well-separated which occurs in junctions with long S regions.
@@ 4.2 Anomalous Proximity Effect for Short S Regions
To further understand the nature of the ZEP and the features observed in the LDOS for the short superconducting segment, we now turn our attention to the the proximity-induced spin-triplet and spin-singlet pair correlations. In Fig. 5, we show the absolute value of the spin-triplet pair correlations $|\vec{d}|$ at the leftmost site of the N region ($i=1$) as a function of energy $E$ and disorder strength $W$ for different lengths of the DN region ($L_{\rm DN}$). In the trivial phase (top row of Fig. 5), we observe a ZEP in $|\vec{d}|$ in the clean limit. As disorder is introduced, the peak splits rapidly, evolving into a fan-shaped structure similar to that observed in the LDOS. Taking line cuts at different values of $W$, we clearly see the flattening of the ZEP into almost a plateau as a function of energy. It is important to clarify why these odd-frequency s-wave triplet pairing, normally robust to scalar disorder, are fragile here. Their fragility is inherited from their parent state; since the trivial ABS is not topologically protected and is easily gapped out by disorder (as we discussed in Sec.4.1), the pair correlations it induces are consequently suppressed at zero energy. Instead, we have a dip at zero energy with two peaks at finite energy corresponding to the energy aquired by the ABS. Note that, however, due to disorder-averaging, the two peaks are smeared out and become the plateau structure observed here. The spatial profile of the spin-triplet pairing at $E\approx 0$ (Fig. 5(e)) shows that in the clean limit (grey curve), triplet correlations exhibit a large value at the DN/S interface where the ABS is localized and it penetrates significantly into the N region. However, with increasing disorder (red and cyan curves), the magnitude of these zero-energy triplet correlations is substantially reduced throughout the entire N region, and their characteristic oscillatory pattern is dampened, especially within the DN segment.
@@ Acknowledgements.
We thank S. Ikegaya and N. Shannon for fruitful discussions. E. A. acknowledges financial support from Nagoya University and Mitsubishi Foundation. Y. T. acknowledges support from JSPS with Grants-in-Aid for Scientific Research (KAKENHI Grant Nos. 23K17668,24K00583,24K00578,24K00556,25H00609, and 25H00613). J. C. acknowledges financial support from the Carl Trygger’s Foundation (Grant No. 22: 2093), the Sweden-Japan Foundation (Grant No. BA24-0003), and the Swedish Research Council (Vetenskapsrådet Grant No. 2021-04121).